OpenStudy (anonymous):

If one measurement of a golden rectangle is 5.9 inches, which number could be the other measurement?.... A) 7.518 ... B) 4.282 ... C) 3.646 ... D) 1.618

OpenStudy (aravindg):

do you know what a golden rectangle is ?

OpenStudy (aravindg):

@Ruhhr ?

OpenStudy (anonymous):

It's l/s = s/l-s.. I know it's suppose to be around 1.618. I'm just confused on how to do this problem

OpenStudy (aravindg):

A golden rectangle is one whose side lengths are in the golden ratio, \[\large 1:\dfrac{1+\sqrt{5}}{2}\]

OpenStudy (anonymous):

Okay so, how woul I do this problem?

OpenStudy (aravindg):

now can you try i t ?

OpenStudy (aravindg):

just use the ratio !!

OpenStudy (aravindg):

@Ruhhr still confused ?

OpenStudy (anonymous):

Yes.. I am still confused..

OpenStudy (aravindg):

well there are only 2 possibilities: either the given side is the side having 1 part in the ratio or it is the the side having part phi

OpenStudy (aravindg):

check both :i will show u hw to do first one :

OpenStudy (aravindg):

so if the given side is having part 1..then other side will be phi times this side ie \[1.61 \times 5.9=9.499\]

OpenStudy (aravindg):

so it is not definitely the first case

OpenStudy (aravindg):

try the second case @Ruhhr

OpenStudy (anonymous):

I got help from a teacher in the school. I take an online class so it's a ltle confusing to me, it was c